OpenStudy (anonymous):

Use differentials to approximate the maximum error propagated when calculating the volume of a sphere, if radius is 5 +/- .01 meters.

OpenStudy (anonymous):

@Nishant_Garg

OpenStudy (anonymous):

Ok so first of all we need to formulate our function, volume of a sphere is a function of it's radius. Thus we define \[V(r)=\frac{4}{3}\pi r^3\] And the formula for differentials as we have it is \[dy=\frac{dy}{dx}.dx\] Or in our case, \[dV=\frac{dV}{dr}.dr\] When a change is made in radius \[\Delta r\] a corresponding change in volume is then \[\Delta V\] When these changes are small, not infinitely small but still quite small, we can approximate for the differentials \[dr \approx \Delta r \space \space \space dV \approx \Delta V\] \[\Delta V \approx \frac{dV}{dr}.\Delta r\] Now we are given the error in the radius, which is our delta r, it is the small change that may occur in the radius \[\Delta r=\pm 0.1\] Thus now we can easily find delta V or the error expected in volume \[\Delta V=\pm \frac{dV}{dr} \times 0.1\] and of course we will put the actual value of r when we compute the derivative so infact I should say\[\Delta V=\pm f'(5) \times 0.1\]

OpenStudy (anonymous):

First find\[\frac{dV}{dr}\] then put the value of r, which is given as 5 in the derivative, can u do that :)

OpenStudy (anonymous):

I though radius is 5 +/- .01 so we will have 2 different radius for this problem right

OpenStudy (anonymous):

No, our radius is only one and that is 5 meters, the plus minus .01 is actually the error in the radius, what we are trying to say is the radius may be more or less by 0.01 units, similarly using the plus minus notation for resultant error in volume we are trying to say that the volume may be more or less than the original amount by the value we've calculated

OpenStudy (anonymous):

sorry I should've actually written \[\Delta V=\pm f'(5) \times 0.01\]

OpenStudy (anonymous):

hmm okay makes sense now. Alright so dv/dr means the direvitive of the volume/ the radius?

OpenStudy (anonymous):

Yep! derivative of the volume function with respect to the radius

OpenStudy (anonymous):

4pi r^2/ 5

OpenStudy (anonymous):

yep! 4pi r^2, now just plug in r=5

OpenStudy (anonymous):

Leave pi as it is, don't use any value for it

OpenStudy (anonymous):

cool! so 100pi/5

OpenStudy (anonymous):

There's no 5 in the denominator! \[\frac{d}{dr}(\frac{4}{3}\pi r^3)=\frac{4}{3}\pi \frac{d}{dr}(r^3)=\frac{4\pi}{3}.3 r^2=4\pi r^2\]

OpenStudy (anonymous):

so 100pi, now we just plug that into \[\Delta V=\pm f'(5) \times 0.01\] in place of f'(5)

OpenStudy (anonymous):

plus minus 100pi *.01

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

now do u want me to go ahead and substitute the value for pi?

OpenStudy (anonymous):

don't put the value of pi yet, rather I'd want to you to simplify that 100 and 0.01

OpenStudy (anonymous):

+/- 1

OpenStudy (anonymous):

+/- 1 pi

OpenStudy (anonymous):

yep! \[\Delta V=\pm \pi\] Now if you want you may use 3.1416 or whatever value of pi you want to

OpenStudy (anonymous):

or u can leave it at that

OpenStudy (anonymous):

I will leave it like that. :)

OpenStudy (anonymous):

So if the radius is more or less by 0.01 than the original value, the volume would be more or less by about 3.1416 cubic metres!

OpenStudy (anonymous):

yes, tysm makes sense now. I was thinking two radius haha was on the wrong track

OpenStudy (anonymous):

It's alright, take care!

OpenStudy (anonymous):

I have another question. it's about Newton's method I need serious help with that one. do u know that? I can open that as a new question

OpenStudy (anonymous):

hmm, not really familiar with Newton's method, I've heard about it but never encountered in syllabus anywhere so I didn't bother either, maybe ask someone else?It'd be worse if I were to misdirect you accidently!

OpenStudy (anonymous):

hm okay let me openanother question than. U might know that one.

OpenStudy (anonymous):

sure